OpenStudy (anonymous):

How to find all the lines through the point (2,1) that are tangent to the parabola y=x^2?

jimthompson5910 (jim_thompson5910):

if y = x^2, then dy/dx = 2x So any tangent line on y = x^2 has a slope of 2x, assuming you're at the point (x,y), where y = x^2

jimthompson5910 (jim_thompson5910):

This means that the slope is m = 2x This line must go through (2,1), so that point helps us to uniquely define the line or lines

jimthompson5910 (jim_thompson5910):

Any point on y = x^2 is of the form (x,y) which is really (x, x^2) since y = x^2 The tangent line must go through the two points (x, x^2) and (2,1)

jimthompson5910 (jim_thompson5910):

The tangent line has the slope m = 2x so if we can connect the two points, and the slope, then we can solve for x

jimthompson5910 (jim_thompson5910):

turns out we can since we can use the slope formula m = (y2 - y1)/(x2 - x1) m = (1 - x^2)/(2 - x) 2x = (1 - x^2)/(2 - x) 2x(2-x) = 1 - x^2 4x - 2x^2 = 1 - x^2 4x - 2x^2 - 1 + x^2 = 0 -x^2 + 4x - 1 = 0 x^2 - 4x + 1 = 0 Now use the quadratic formula to solve for x. I'll let you do this. Tell me what you get.